• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 16:40:39 on March 7,2005

    Entry number 140898

    This entry is a followup to: 140888


    keyword=ctimel (Left) problems again

    Olivier reports that L-arm TC ("ctimel") shifted again 80 nsec.
    So I put it back to normal. Then Olivier took a run and saw TC was
    bouncing around by +/- 80 nsec. It looks like the CAMAC delay module
    that is used to make this delay is flakey. Tomorrow when we have
    access (hall closes at 5 pm today) I'll re-route this signal to avoid
    the presumably bad module. (Note, 80 nsec is a multiple of the LSB.)